OpenStudy (anonymous):

A bag contains 2 steel balls and 5 brass balls. The total weight is 13 pounds. If 2 steel balls are added and 2 brass balls are removed, the weight decreases to 12 pounds how much does each kind of ball weigh? A. Steel: 4 pounds;brass:1pound B. steel:1.5 pounds;brass:2 pounds C.steel:0.75 pounds;brass:3pounds D.steel:3 pounds;brass:4 pounds

OpenStudy (anonymous):

Let \(s\) and \(b\) represent the weights of the steel and brass balls, respectively. Then you have \[\begin{cases} 2s+5b=13\\ 4s+3b=12 \end{cases}\] Initially, you have 2 steel and 5 brass which weigh 13 pounds. Then you add 2 steel and take out 2 brass, and the weight drops by a pound. Solve the system.